Average of 5 numbers is 80.If three of these are 70,80 and 90.Find the average of last two?
Answer by Cromlix(4381)

Average of 5 numbers is 80
Total of 5 numbers = 80 * 5 = 400
70 + 80 + 90 = 240
400 - 240
=160
Average of the last two numbers is;
160/2 = 80.
Hope this helps.
:-)